# Sharp concentration inequality for the sum of random variables

### v1

@@ 3 Classical sum concentration profiles
In practice, the tail quantile operator $T_{X}$ is rarely available in closed form. It is therefore natural to work instead with tractable envelopes for the survival operator $S_{\mu}$, and to propagate such bounds to $S_{X_{1}+\cdots+X_{n}}$ through Corollary 1.
For simplicity, for any $a>0$, let us introduce the operator $\mathrm{Id}^{-a}:\mathbb{R}\to 2^{\mathbb{R}}$ defined by
$\mathrm{Id}^{-a}(t)=\{t^{-a}\}\ \text{for }t>0,\qquad\mathrm{Id}^{-a}(t)=\emptyset\ \text{for }t\leq 0.$
Elementary integrations yield the following Hardy-transform identities:
@@
Inverting these identities (the order is preserved since all the mappings involved are nonincreasing), and applying Corollary 1, one obtains the following explicit profiles: for any constant $C>0$,
@@
  • if $S_{\mu}\leq C\,\mathcal{E}_{1}$ where the notation $\mathcal{E}_{1}:t\mapsto\{e^{-t}\}$ was taken from Louart (2024), then
@@
When regular inequalities like $S_{\mu}\leq C\,\mathrm{Id}^{-q}$ and $S_{\mu}\leq C\,\mathcal{E}_{1}$ are not satisfying or even reachable, it is worth trying to compare $S_{\mu}$ to $\mathrm{Id}^{-q}$ or $\mathcal{E}_{1}$ in the convex transformation order van Zwet (1964) to obtain simple and sharp bounds as shown in Corollary 1. We now prove this result.
Proof of Corollary 1.
@@
[…]
@@
$\mathcal{H}(T_{\mu})(p)\leq\mathcal{H}(\alpha^{-1})(p)\leq\bigl(\alpha^{-1}\circ\mathrm{Id}^{-q}\bigr)\!\left(\frac{q}{q-1}p^{-\frac{1}{q}}\right)=\alpha^{-1}\!\left(\left(\frac{q-1}{q}\right)^{q}p\right).$
Combining this with Theorem 1 in the i.i.d. marginal case gives
$T_{\frac{1}{n}\sum_{k=1}^{n}X_{k}}(p)\leq\mathcal{H}(T_{\mu})(p)\leq\alpha^{-1}\!\left(\left(\frac{q-1}{q}\right)^{q}p\right).$
@@
$S_{X_{1}+\cdots+X_{n}}\leq\left(\frac{q}{q-1}\right)^{q}\,\alpha,$
which is the first result. The second result is shown the same way relying on the concavity of $\alpha^{-1}\circ\mathcal{E}_{1}$ and (13). ∎
@@
one may wonder whether the two convexity assumptions (power-type versus exponential-type) overlap. In general they do not; the next theorem shows that they coincide exactly in the limit $q\to\infty$.
Theorem 3.
Given a continuous mapping $f:(0,\infty)\to\mathbb{R}$, if for every $q>0$ the function $f\circ\mathrm{Id}^{-q}$ is convex (resp. concave), then the function $f\circ\mathcal{E}_{1}$ is convex (resp. concave). If we assume in addition that $f$ is nondecreasing (resp. nonincreasing), then the converse is true.
The theorem relies on the following limiting representation of the geometric mean.
Lemma 6 (Power–geometric limit).
@@
Proof.
Set $r:=1/q$, $\alpha:=1/a$ and $\beta:=1/b$. Then
$m_{1/r,\lambda}(a,b)=\bigl(\lambda\alpha^{r}+(1-\lambda)\beta^{r}\bigr)^{-1/r},\qquad\text{so}\qquad-\log\bigl(m_{1/r,\lambda}(a,b)\bigr)=\frac{1}{r}\log\bigl(\lambda\alpha^{r}+(1-\lambda)\beta^{r}\bigr).$
@@
$\frac{1}{r}\Bigl(\log g(r)-\log g(0)\Bigr)=\frac{1}{r}\int_{0}^{r}\frac{g^{\prime}(s)}{g(s)}\,ds,\qquad g(s):=\lambda\alpha^{s}+(1-\lambda)\beta^{s},$
(which appears for instance in Qi et al. (2000)), we obtain
$-\log\bigl(m_{1/r,\lambda}(a,b)\bigr)$ $=\frac{1}{r}\int_{0}^{r}\frac{\lambda\alpha^{s}\log(\alpha)+(1-\lambda)\beta^{s}\log(\beta)}{\lambda\alpha^{s}+(1-\lambda)\beta^{s}}\,ds$
$\underset{r\to 0}{\longrightarrow}\lambda\log(\alpha)+(1-\lambda)\log(\beta)=-\lambda\log(a)-(1-\lambda)\log(b),$
by continuity of the integrand. Taking exponentials yields the claim. ∎
Proof of Theorem 3.
@@
Letting $q\to\infty$ and using Lemma 6 together with continuity of $f$, we obtain
@@
Writing $a=e^{-\alpha}$ and $b=e^{-\beta}$, this inequality is exactly the convexity of $f\circ\mathcal{E}_{1}$.
Conversely, assume that $f$ is nondecreasing and that $f\circ\mathcal{E}_{1}$ is convex. Fix $q>0$ and $a,b>0$, and write $a=e^{\alpha}$, $b=e^{\beta}$ with $\alpha=\log(a)$ and $\beta=\log(b)$. Since $\exp$ is convex, we have
$\lambda e^{\alpha}+(1-\lambda)e^{\beta}\geq e^{\lambda\alpha+(1-\lambda)\beta}.$
@@
$f\bigl((\lambda a+(1-\lambda)b)^{-q}\bigr)$ $=f\bigl((\lambda e^{\alpha}+(1-\lambda)e^{\beta})^{-q}\bigr)$
$\leq f\bigl(e^{-q(\lambda\alpha+(1-\lambda)\beta)}\bigr).$
@@
$f\bigl(e^{-q(\lambda\alpha+(1-\lambda)\beta)}\bigr)$ $\leq\lambda f(e^{-q\alpha})+(1-\lambda)f(e^{-q\beta})=\lambda f(a^{-q})+(1-\lambda)f(b^{-q}),$
which proves that $f\circ\mathrm{Id}^{-q}$ is convex. ∎

### v2

@@ 3 Classical sum concentration profiles
In practice, the tail quantile operator $T_{X}$ is rarely available in closed form. It is therefore natural to work instead with tractable envelopes for the survival operator $S_{\mu}$, and to propagate such bounds to $S_{\frac{1}{n}\sum_{k=1}^{n}X_{k}}$ through Corollary 1.5.
Elementary calculations yield the following Hardy-transform identities (with the power-operator notation introduced in (5)):
@@
Inverting these identities (inversion preserves the order since all the mappings involved are nonincreasing), and applying Corollary 1.5, one obtains the following explicit profiles: for any constant $C>0$,
@@
  • if $S_{\mu}\leq C\,\mathcal{E}_{1}$, where the notation $\mathcal{E}_{1}:t\mapsto\{e^{-t}\}$ is taken from [Louart(2024)], then
@@
When simple pointwise bounds such as $S_{\mu}\leq C\,\mathrm{Id}^{-q}$ or $S_{\mu}\leq C\,\mathcal{E}_{1}$ are not satisfactory, or are not available, it is worth comparing $S_{\mu}$ to $\mathrm{Id}^{-q}$ or $\mathcal{E}_{1}$ with respect to the convex transformation order [van Zwet(1964)]. This yields simple and sharp bounds, as stated in Corollary 1.5, which we now prove.
Proof 3.1 (Proof of Corollary 1.5).
@@
Now assume that for all $i\in[n]$$S_{X_{i}}\leq\alpha$. By (1), this implies $T_{X_{i}}\leq\alpha^{-1}$, and since the Hardy transform is order-preserving, we obtain $\mathcal{H}(T_{X_{i}})\leq\mathcal{H}(\alpha^{-1})$. For $p>0$,
@@
$\mathcal{H}\left(\frac{1}{n}\sum_{k=1}^{n}T_{X_{k}}\right)(p)\leq\mathcal{H}(\alpha^{-1})(p)\leq\bigl(\alpha^{-1}\circ\mathrm{Id}^{-q}\bigr)\!\left(\frac{q}{q-1}p^{-\frac{1}{q}}\right)=\alpha^{-1}\left(\left(\frac{q-1}{q}\right)^{q}p\right).$
Combining this with Theorem 1.2 in the identically distributed case gives
$T_{\frac{1}{n}\sum_{k=1}^{n}X_{k}}(p)\leq\mathcal{H}\left(\frac{1}{n}\sum_{k=1}^{n}T_{X_{k}}\right)(p)\leq\alpha^{-1}\!\left(\left(\frac{q-1}{q}\right)^{q}p\right).$
@@
$S_{\frac{1}{n}\sum_{k=1}^{n}X_{k}}\leq\left(\frac{q}{q-1}\right)^{q}\,\alpha,$
which is the first result.
The second result is proved analogously, using the concavity of $\alpha^{-1}\circ\mathcal{E}_{1}$ (equivalently, the convexity of $\log\circ\alpha$) and the identity $\mathcal{H}(-\log)=1-\log$ from (15).
@@
one may wonder whether the two convexity assumptions (power-type versus exponential-type) overlap. In general they do not; the next proposition shows that they coincide exactly in the limit $q\to\infty$.
Proposition 3.2.
Given a continuous function $f:(0,\infty)\to\mathbb{R}$, if for every $q>0$ the function $f\circ\mathrm{Id}^{-q}$ is convex (resp. concave), then the function $f\circ\mathcal{E}_{1}$ is convex (resp. concave). If we assume in addition that $f$ is nondecreasing (resp. nonincreasing), then the converse is true.
The proposition relies on the following limiting representation of the geometric mean.
Lemma 3.3 (Power–geometric limit).
@@
Proof 3.4.
Set $r:=1/q$, $A:=1/a$ and $B:=1/b$. Then
$m_{1/r,\lambda}(a,b)=\bigl(\lambda A^{r}+(1-\lambda)B^{r}\bigr)^{-1/r},\qquad\text{so}\qquad-\log\bigl(m_{1/r,\lambda}(a,b)\bigr)=\frac{1}{r}\log\bigl(\lambda A^{r}+(1-\lambda)B^{r}\bigr).$
@@
$\frac{1}{r}\Bigl(\log g(r)-\log g(0)\Bigr)=\frac{1}{r}\int_{0}^{r}\frac{g^{\prime}(s)}{g(s)}\,ds,\qquad g(s):=\lambda A^{s}+(1-\lambda)B^{s},$
(which appears for instance in [Qi et al.(2000)Qi, Mei, Xia, and Xu]), we obtain
$-\log\bigl(m_{1/r,\lambda}(a,b)\bigr)$ $=\frac{1}{r}\int_{0}^{r}\frac{\lambda A^{s}\log(A)+(1-\lambda)B^{s}\log(B)}{\lambda A^{s}+(1-\lambda)B^{s}}\,ds$
$\underset{r\to 0}{\longrightarrow}\lambda\log(A)+(1-\lambda)\log(B)=-\lambda\log(a)-(1-\lambda)\log(b),$
by continuity of the integrand. Taking exponentials yields the claim.
Proof 3.5 (Proof of Proposition 3.2).
@@
Letting $q\to\infty$ and using Lemma 3.3 together with continuity of $f$, we obtain
@@
Writing $a=e^{-x}=\mathcal{E}_{1}(x)$ and $b=\mathcal{E}_{1}(y)$, $a^{\lambda}b^{1-\lambda}=\mathcal{E}_{1}(\lambda x+(1-\lambda)y)$ and this inequality is exactly the convexity of $f\circ\mathcal{E}_{1}$.
Conversely, assume that $f$ is nondecreasing and that $f\circ\mathcal{E}_{1}$ is convex. Fix $q>0$ and $a,b>0$, and write $a=e^{x}$, $b=e^{y}$ with $x=\log(a)$ and $y=\log(b)$. Since $\exp$ is convex, we have
$\lambda e^{x}+(1-\lambda)e^{y}\geq e^{\lambda x+(1-\lambda)y}.$
@@
$f\bigl((\lambda a+(1-\lambda)b)^{-q}\bigr)$ $=f\bigl((\lambda e^{x}+(1-\lambda)e^{y})^{-q}\bigr)$
$\leq f\bigl(e^{-q(\lambda x+(1-\lambda)y)}\bigr).$
@@
$f\bigl(e^{-q(\lambda x+(1-\lambda)y)}\bigr)$ $\leq\lambda f(e^{-qx})+(1-\lambda)f(e^{-qy})=\lambda f(a^{-q})+(1-\lambda)f(b^{-q}),$
which proves that $f\circ\mathrm{Id}^{-q}$ is convex.